OpenStudy (anonymous):

5 1/2% of $568.85 A.$31.29 B.312.87

OpenStudy (anonymous):

my answer is $31.29 is it correct?

OpenStudy (anonymous):

yah it is

OpenStudy (surryyy):

5 1/2% of $568.85 = 31.28675 31.28675 rounded to the nearest hundredth = 31.29

OpenStudy (jhannybean):

Hah, nevermind. I got it :)

OpenStudy (surryyy):

lol. u made me confused

OpenStudy (jhannybean):

\[\frac{\frac{11}{2}}{100} = \frac{11}{200}\cdot 568.85 \approx 31..29\]

OpenStudy (jhannybean):

I was confused by the notation :P

OpenStudy (surryyy):

lol

OpenStudy (jhannybean):

\[5\frac{1}{2}\% = \frac{11}{2}\% = \frac{11}{200}\]